# Semi-orthogonality in Fukaya-Seidel mirrors to blowups of abelian varieties

### v1

@@ 3.2.1. Wrapping conventions in the cochain complex
If they are not transverse we perturb by isotopies. Let $L$ be a Lagrangian of the form $L_{j}$ or $K_{j}$. Recall that a Lagrangian isotopy is a homotopy
@@
which is a smooth embedding. We have a horizontal distribution
@@
for the symplectic form $\omega$ because $W_{0}$ is a symplectic fibration, that is, $\omega$ restricts to a symplectic form on smooth fibers. Isotopies in a Landau-Ginzburg model [ACLL23b, §2] are constructed by parallel transport over homotopies $\delta(s,t)$ of curves in the base $\mathbb{C}$. Here they partially rotate the angular ends going to infinity, not crossing $\pi$, and fully rotate around 0 in the other direction.
Let $\rho^{t}$ be the flow of the horizontal lift of the Hamiltonian vector field for the Hamiltonian $H$ defined as follows. Let $w$ be the coordinate on the base $\mathbb{C}^{*}$ of $W_{0}$. For $|w|\neq 0$ and $\chi(\arg w):[-\pi,\pi)\to[0,1]$ a bump function supported away from $\pi$, pick $0<R_{1}<R_{2}<T^{\epsilon}<S<R_{3}<R_{4}<T^{\varepsilon}$ and define
(3.3) $H(w):=\begin{cases}\chi(\arg w)\cdot|w|&R_{4}<|w|<T^{\varepsilon}\\ \text{interpolate}&|w|\in[R_{3},R_{4}]\\ 0&R_{2}<|w|<R_{3}\\ \text{interpolate}&|w|\in[R_{1},R_{2}]\\ |w|^{2}&|w|<R_{1}\end{cases}$
where recall $|W_{0}|:Y_{0}\to[0,T^{\varepsilon}]$ arises as $W_{0}$ is a fibration only over a small disc. So within this domain but outside of the annulus
@@
Lagrangians project under $W_{0}$ to arcs which wrap slightly on one end and fully around 0. Within the smaller annulus $A=\{w\in\mathbb{C}\mid R_{2}<|w|<R_{3}\}$ there is no movement, including at S-S where Lagrangians are defined as parallel transport from in Equation (3.1). Define
$\rho^{t}:Y^{0}\to Y^{0}$
to be the flow of a vector field $\xi^{\#}$ which is the horizontal lift of $\xi$ defined by $\iota_{\xi}\omega_{\mathbb{C}}=dH$ for $\omega_{\mathbb{C}}=\frac{i}{2}dw\wedge d\overline{w}=dw_{x}\wedge dw_{y}=rdr\wedge d\theta=e^{2u}du\wedge d\theta$ for $w=w_{x}+iw_{y}=re^{i\theta}=e^{u+i\theta}$. (Note that $\xi$ is sometimes called $X_{H}$ in literature where $\iota_{X_{H}}\omega_{\mathbb{C}}=-dH$, but in [AA24] $X_{H}$ is used for a wrapping Hamiltonian $H$ in the fiber while $\xi$ realizes the wrapping in the base. Since we are wrapping in the base and not the fiber, to be consistent with them, we use $\xi$.) By [ACLL23b, Remark 3.2], $\omega_{\mathbb{C}}$ is related to $\omega$ by
The relation between positively or negatively wrapping with respect to $\lambda:=\frac{1}{2}r^{2}d\theta$ (where $\omega_{\mathbb{C}}=d\lambda$) and counterclockwise/clockwise wrapping around 0 and infinity is as follows.
@@
    (3.6) $\iota_{-\chi^{\prime}\partial_{r}+\frac{\chi}{r}\partial_{\theta}}(rdr\wedge d\theta)=-dH=-d(\chi(\theta)r)$
    so on $\xi|_{\{R_{4}<r<T^{\varepsilon}\}}=-\chi^{\prime}\partial_{r}+\frac{\chi}{r}\partial_{\theta}$ we find that
    (3.7) $\lambda(-\chi^{\prime}\partial_{r}+\frac{\chi}{r}\partial_{\theta})=\frac{1}{2}r^{2}d\theta(\frac{\chi}{r}\partial_{\theta})=\frac{1}{2}\chi r>0$
@@
    Around 0 we flow clockwise by $\xi=2\frac{\partial}{\partial\theta}$ as we approach 0. This is because under the conformal coordinate change $w\mapsto\tilde{w}:=1/w$ on $\mathbb{C}^{*}$, then
    (3.8) $\omega_{\mathbb{C}}=\frac{1}{\tilde{r}^{3}}d\tilde{r}\wedge d\tilde{\theta}$
    and for $\tilde{r}>1/R_{1}$
    (3.9) $dH=\frac{-2}{\tilde{r}^{3}}d\tilde{r}=\iota_{-2\partial/\partial\tilde{\theta}}\frac{1}{\tilde{r}^{3}}d\tilde{r}\wedge d\tilde{\theta}$
@@
    (3.10) $\lambda(-2\frac{\partial}{\partial\tilde{\theta}})=-\frac{1}{2\tilde{r}^{2}}d\tilde{\theta}(-2\frac{\partial}{\partial\tilde{\theta}})=\frac{1}{\tilde{r}^{2}}>0$
    so we wrap counterlockwise with increasing $\tilde{r}$. Inverting to $w$, we wrap clockwise with decreasing $r$.
Now let $L^{k}:=\rho^{-\tilde{\epsilon}k}(L)$ for small positive number $\tilde{\epsilon}$. So larger $k$ means more negatively wrapped. Following [AA24, Equation (3.40)], we formally define a directed category $\mathcal{O}$ with objects the Lagrangians $\{L^{j}\}_{j\in\mathbb{Z}}$ so they form an exceptional collection
@@
The differential $\partial:CF^{m}(L_{0},L_{1})\to CF^{m+1}(L_{0},L_{1})$ on the cochain complex of transversely intersecting Lagrangians is
(3.11) $\partial(p)=\sum_{q\in L_{0}\cap L_{1}}\sum_{\begin{subarray}{c}[u]:[u]=\beta\in H_{2}(Y^{0},L_{0}\cup L_{1})\\ \text{ind}(u)=1\end{subarray}}\#\mathcal{M}(p,q;[u],J)/(u\sim u_{a})e^{-2\pi\int_{\beta}\omega}q$
@@
(3.12) $\mathcal{M}(p,q;[u],J):=\left\{u:(\mathbb{R}\times[0,1],\mathbb{R}\times\{0\}\cup\mathbb{R}\times\{1\})\to(Y^{0},L_{0}\cup L_{1})\middle|\begin{array}[]{l}\overline{\partial}_{J}(u)=0,\ E(u)<\infty,\\ u(\mathbb{R},0)\subset L_{0},\ u(\mathbb{R},1)\subset L_{1},\\ \lim_{s\to-\infty}u(s,t)=q,\\ \ \lim_{s\to\infty}u(s,t)=p\end{array}\right\}$
where we quotient by $u_{a}(s,t)=u(s-a,t)$ for each $a\in\mathbb{R}$, the translation in the $s$ direction, and where $J$ is an almost complex structure compatible with $\omega$ which is also regular, meaning the linearization of $\overline{\partial}_{J}$ is surjective for all $u\in\mathcal{M}(p,q;[u],J)$. In order for this to be well-defined, meaning in order for us to be able to count the moduli space, it suffices for it to be a 0-dimensional compact smooth manifold. The 0-dimensional part is ensured by choosing $p$ and $q$ so that $\deg(p)+1=\deg(q)$. The compact part is ensured by the curves $u$ satisfying a maximum principle, the way holomorphic curves do, and by excluding bubbling. And the smooth manifold part can be ensured by the existence of regular $J$.
This was proven in [Can20] in the setting for $(Y,-v_{0})$. The difference between the set-up in [Can20] and here is that we have a puncture at 0 in the base so therefore we have additional Lagrangians fibered over single arcs that can wind infinitely many times around 0 in the base. So we have several cases to consider.
@@ 3.2.2. Algorithm to derive the category
The morphisms are given by the wrapped Lagrangian Floer cohomology theory. They are defined by the property that Hamiltonian-isotopic Lagrangians should be quasi-isomorphic; they should be isomorphic when we derive the category. There is a convenient tool from homotopy theory that does the job, called a homotopy colimit [AS10, AA24]. As explained in the introduction to [Hir03], a goal of homotopy theory is to identify homotopic maps, but a homotopy on topological spaces may not make sense in other categories. To identify homotopic morphisms $f$ and $g$, we define a set called the weak equivalences to be where fgf-g lands in, and then localize the category at weak equivalences. To take a homotopy colimit of a diagram one needs a model category: a category with three sets of morphisms called weak equivalences, fibrations, and cofibrations. Quillen axiomatized such a category as a place where one can do homotopy theory, calling them “a category of models for a homotopy theory” [Qui67]. Note that more modern approaches in homotopy theory use $\infty$-categories instead of model categories. In the wrapped Fukaya category, weak equivalences will be multiplication by quasi-units. Here is the algorithm used to invert the quasi-units:
@@
    For chain complexes bounded below, over an abelian category with enough projectives, there is a model category structure where the weak equivalences are the quasi-isomorphisms, the fibrations are the epimorphisms, and the cofibrations are injective maps $i:C_{\bullet}\to D_{\bullet}$} where the cokernel complex ${\text{Coker}}(i)_{\bullet}$} is a complex of projective objects [Qui67, §1.2 Example B]. It follows that the cofibrant objects are the complexes whose objects are all projective. In our case ([AA24, Equation (3.40)], [Can20, Equation (4.2)]) the model category will be a category of modules over another category $\mathcal{O}$,
@@
    where id is the identity morphism when composed with any $CF^{*}(L_{0}^{k_{0}},L_{1}^{k_{1}})$. Then the model category $\mathcal{A}$ is the category of $\mathcal{O}$-modules with objects $\mathcal{Y}_{L}$ given by Yoneda modules $X\mapsto\mathcal{O}(X,L)$ and quasi-isomorphisms are the quasi-units. Counts of curves between multiple intersection points define the $A_{\infty}$-structure $\mu^{k}$.
@@
    Construct the homotopy category by deriving $\mathcal{A}$ so that quasi-isomorphisms become isomorphisms. An object is therefore an equivalence class of a chain complex, up to quasi-isomorphism, and a morphism is an equivalence of a chain map, up to chain homotopy. This is a localization process [Wei94, §10.3] which has an $A_{\infty}$-version in [LO06].
@@
      The directed category $\mathbb{Z}$ indexes the diagram for each Lagrangian $L$, given by $\ldots\to\mathcal{Y}_{L^{k}}\xrightarrow[]{\mu^{2}(e_{L^{k}},-)}\mathcal{Y}_{L^{k+1}}\to\ldots$ for $k\in\mathbb{Z}$ given by multiplication by quasi-units. Define
      $\mathcal{Y}_{L^{\infty}}:=\text{hocolim}(\mathcal{Y}_{L^{k}}).$
@@
The homotopy direct limit over a collection of isotopies avoids the definition depending on a given isotopy. Note that using a quadratic Hamiltonian to wrap as in [Aur14, §4.2] is equivalent to a colimit by linear Hamiltonians (the original definition of the wrapped Fukaya category [AS10, §5.1]) by [OT24, Proposition 6.1.9].
@@ 3.2.3. Quasi-isomorphisms: quasi-units
Following the convention of [AA24, Equation (3.42)], a quasi-unit counts certain curves $\Sigma:=\mathbb{D}\backslash\{1\}\to Y^{0}$ with one moving Lagrangian boundary condition $\rho^{-\tilde{\epsilon}g_{e}(s)}(L)$ for $g_{e}:\mathbb{R}\cong\partial\Sigma\to[k,k+1]$ monotonically increasing smooth and constant near the ends, meaning there is some $K\in\mathbb{R}$ so
@@
so the Lagrangian isotopy is traversed backwards along the boundary.
For the PSS isomorphism $PSS:H^{*}(L)\to HF^{*}(L^{k},L^{k+1})$, the quasi-unit is the image of 1, and can be constructed on the chain-level by a disc count,
(3.15) $e_{L^{k}}:=PSS(1)=\left[\sum_{p\in L^{k}\cap L^{k+1}}\sum_{\begin{subarray}{c}[u]:[u]=\beta\in H_{2}(Y^{0},\bigcup_{s\in\mathbb{R}}\rho^{-\tilde{\epsilon}g_{e}(s)}(L^{k}))\\ \text{ind}(u)=0\end{subarray}}\#\mathcal{M}(p;[u],J)e^{-2\pi\int_{\beta}\omega}\cdot p\in CF(L^{k},L^{k+1})\right]\\$

### v2

@@ 3.2.1. Wrapping conventions and the cochain complex
If they are not transverse, we perturb by positive and negative isotopies. We describe what we mean by positive. Let $L$ be a Lagrangian of the form $L_{j}$ or $K_{j}$. Recall that a Lagrangian isotopy is a homotopy
@@
which is a smooth embedding. We have a horizontal distribution, where $Y^{0}_{b}$ denotes the fiber $W_{0}^{-1}(b)$, given by
@@
for the symplectic form $\omega$ because $W_{0}$ is a symplectic fibration, that is, $\omega$ restricts to a symplectic form on smooth fibers. Isotopies in a Landau-Ginzburg model [ACLL_LG, §2] are constructed by parallel transport over homotopies $\delta(s,t)$ of curves in the base $\mathbb{C}$.
Let $\rho^{t}$ be the flow of the horizontal lift of the Hamiltonian vector field for the Hamiltonian $H$ defined as follows. Let $w$ be the complex coordinate on the base $\mathbb{C}^{*}$ of $W_{0}$. There are several commonly used real coordinate systems on the base. We express them here for completeness. We have
(3.6) $w=w_{x}+iw_{y}=re^{2\pi i\theta}=e^{2\pi(\mu+i\theta)}$
where the last expression is in action-angle coordinates $(\mu,\theta)$ for $\mu(r)=\frac{1}{2\pi}\log r$. By [ACLL_LG, Remark 3.2], the standard form $\omega_{\mathbb{C}}$ is related to $\omega$ by
Here on the base we have $\mathbb{C}^{*}$ instead of $\mathbb{C}$, so for $w$ the coordinate on $\mathbb{C}^{*}$ and under the map $\exp(2\pi\cdot):\mathbb{C}\to\mathbb{C}^{*}$ the standard form on $\mathbb{C}$ becomes the following on $\mathbb{C}^{*}$
(3.8) $\omega_{\mathbb{C}^{*}}=\frac{i}{2}d\left(\frac{1}{2\pi}\log w\right)\wedge d\left(\frac{1}{2\pi}\log\overline{w}\right)=\frac{i}{2}\frac{dw\wedge d\overline{w}}{(2\pi)^{2}|w|^{2}}=d\mu\wedge d\theta.$
For $\mu\neq 0$ and $\chi(\theta):[-\pi,\pi)\to[0,1]$ a bump function supported away from $\pi$, consider concentric rings in the base around 0 defined by $0<R_{1}<R_{2}<T^{\epsilon}<S<R_{3}<R_{4}<T^{\varepsilon}$ as in Figure 3.
$\mathbb{C}^{*}$
$R_{1}$
$R_{2}$
$R_{3}$
$R_{4}$
$T^{\varepsilon}$
0
$T^{\epsilon}$
S-S
Figure 3. $Y^{0}$ is a fibration over the $T^{\varepsilon}$ disc about 0. U-shape curves $W_{0}(L_{j})$ pass through $b_{S}=-S$. Arcs $W_{0}(K_{j})$ start from, but don’t include, 0. The singular fiber of $W_{0}$ is above $T^{\epsilon}$ marked by an $\times$.
Define
(3.9) $H(w):=\begin{cases}\chi(\theta)\cdot\mu&R_{4}<|w|<T^{\varepsilon}\\ \text{interpolate}&|w|\in[R_{3},R_{4}]\\ 0&R_{2}<|w|<R_{3}\\ \text{interpolate}&|w|\in[R_{1},R_{2}]\\ \mu&|w|<R_{1}\end{cases}$
where recall $|W_{0}|:Y_{0}\to[0,T^{\varepsilon}]$ because $W_{0}$ is a fibration only over a small disc. So within this domain but outside of the annulus
@@
Lagrangians project under $W_{0}$ to arcs which wrap slightly for large $|w|$ and fully around 0. Within the smaller annulus $A=\{w\in\mathbb{C}\mid R_{2}<|w|<R_{3}\}$ there is no movement, including at S-S where Lagrangians are defined as parallel transport from in Equation (3.1). Define
(3.11) $\rho^{t}:Y^{0}\to Y^{0}$
to be the flow of a vector field $\xi^{\#}$ which is the horizontal lift of $\xi$ defined by $\iota_{\xi}\omega_{\mathbb{C}^{*}}=-dH$. (Note that $\xi$ is sometimes called $X_{H}$ in the literature, but in [AA24] $X_{H}$ is used for a wrapping Hamiltonian $H$ in the fiber while $\xi$ realizes the wrapping in the base. Since we are wrapping in the base and not the fiber, to be consistent with them, we use $\xi$.)
The relation between positively or negatively wrapping with respect to $\lambda:=\mu d\theta$ (where $\omega_{\mathbb{C}^{*}}=d\lambda$) and counterclockwise/clockwise wrapping around 0 and infinity is as follows.
@@
    (3.12) $\iota_{-\chi^{\prime}\mu\partial_{\mu}+\chi\partial_{\theta}}(d\mu\wedge d\theta)=-dH=-d(\chi(\theta)\mu)$
    so on $\xi|_{\{R_{4}<r<T^{\varepsilon}\}}=-\chi^{\prime}\mu\partial_{\mu}+\chi\partial_{\theta}$ we find that
    (3.13) $\lambda(-\chi^{\prime}\mu\partial_{\mu}+\chi\partial_{\theta})=\mu d\theta(-\chi^{\prime}\mu\partial_{\mu}+\chi\partial_{\theta})=\chi\mu>0$
@@
    Around 0 we flow clockwise. This is because under the conformal coordinate change $w\mapsto\tilde{w}:=w^{-1}$ on $\mathbb{C}^{*}$, then $(\mu,\theta)\mapsto(\tilde{\mu},\tilde{\theta})=(-\mu,-\theta)$ and
    (3.14) $\omega_{\mathbb{C}^{*}}=d\tilde{\mu}\wedge d\tilde{\theta}$.
    Thus for $\tilde{r}=r^{-1}>1/R_{1}$
    (3.15) $dH=d\tilde{\mu}=\iota_{-\partial/\partial\tilde{\theta}}d\tilde{\mu}\wedge d\tilde{\theta}$
@@
    (3.16) $\lambda\left(-\frac{\partial}{\partial\tilde{\theta}}\right)=\tilde{\mu}d\tilde{\theta}\left(-\frac{\partial}{\partial\tilde{\theta}}\right)=-\tilde{\mu}<0$
    so we wrap negatively/counterclockwise in $\tilde{\theta}$. Inverting back to $w$, we wrap negatively/clockwise in $\theta$.
Now let
(3.17) $L^{k}:=\rho^{-\tilde{\epsilon}k}(L)$
for small positive number $\tilde{\epsilon}$. So larger $k$ means more negatively wrapped. Following [AA24, Equation (3.40)], we define a directed category $\mathcal{O}$ with objects the Lagrangians $\{L^{j}\}_{j\in\mathbb{Z}}$ so they form an exceptional collection
@@
The differential $\partial:CF^{m}(L_{0},L_{1})\to CF^{m+1}(L_{0},L_{1})$ on the cochain complex of transversely intersecting Lagrangians in Equation (3.5) is
(3.18) $\partial(p)=\sum_{q\in L_{0}\cap L_{1}}\sum_{\begin{subarray}{c}[u]:[u]=\beta\in\pi_{2}(Y^{0},L_{0}\cup L_{1})\\ \text{ind}(u)=1\end{subarray}}\#\mathcal{M}(p,q;[u],J)/(u\sim u_{a})e^{-2\pi\int_{\beta}\omega}\cdot q$
@@
(3.19) $\mathcal{M}(p,q;[u],J):=\left\{u:(\mathbb{R}\times[0,1],\mathbb{R}\times\{0\}\cup\mathbb{R}\times\{1\})\to(Y^{0},L_{0}\cup L_{1})\middle|\begin{array}[]{l}\overline{\partial}_{J}(u)=0,\ E(u)<\infty,\\ u(\mathbb{R},0)\subset L_{0},\ u(\mathbb{R},1)\subset L_{1},\\ \lim_{s\to-\infty}u(s,t)=q,\\ \ \lim_{s\to\infty}u(s,t)=p\end{array}\right\}.$
We quotient by $u\sim u_{a}$ where $u_{a}(s,t)=u(s-a,t)$ for each $a\in\mathbb{R}$, the translation in the $s$ direction, and $J$ is an almost complex structure compatible with $\omega$ which is also regular, meaning the linearization of $\overline{\partial}_{J}$ is surjective for all $u\in\mathcal{M}(p,q;[u],J)$. In order for this to be well-defined, meaning in order for us to be able to count the moduli space, it suffices for it to be a 0-dimensional compact smooth manifold. The 0-dimensional part is ensured by choosing $p$ and $q$ so that $\deg(p)+1=\deg(q)$. The compact part is ensured by the curves $u$ satisfying a maximum principle, the way holomorphic curves do, and by excluding bubbling. And the smooth manifold part can be ensured by the existence of regular $J$.
This was proven in [Ca20] in the setting for $(Y,-v_{0})$. The difference between the set-up in [Ca20] and here is that we have a puncture at 0 in the base so therefore we have additional Lagrangians fibered over single arcs that can wind infinitely many times around 0 in the base. So we have several cases to consider.
@@ 3.2.2. Algorithm to derive the category
As outlined at the start of this chapter, here is the algorithm used to invert the quasi-units:
@@
    For chain complexes bounded below, over an abelian category with enough projectives, there is a model category structure where the weak equivalences are the quasi-isomorphisms, the fibrations are the epimorphisms, and the cofibrations are injective maps $i:C_{\bullet}\to D_{\bullet}$} where the cokernel complex ${\text{Coker}}(i)_{\bullet}$} is a complex of projective objects [Qu67, §1.2 Example B]. It follows that the cofibrant objects are the complexes whose objects are all projective. In our case ([AA24, Equation (3.40)], [Ca20, Equation (4.2)]) the model category will be a category of modules over another category $\mathcal{O}$,
@@
    where id acts as the identity morphism when composed with any $CF^{*}(L_{0}^{k_{0}},L_{1}^{k_{1}})$. Then the model category $\mathcal{A}$ is the category of $\mathcal{O}$-modules with objects $\mathcal{Y}_{L}$ given by Yoneda modules $X\mapsto\mathcal{O}(X,L)$ and quasi-isomorphisms are the quasi-units, which are defined in the next subsection, §3.2.3. Counts of curves between multiple intersection points define the $A_{\infty}$-structure $\mu^{k}$.
@@
    Construct the homotopy category by deriving $\mathcal{A}$ so that quasi-isomorphisms become isomorphisms. An object is therefore an equivalence class of a chain complex, up to quasi-isomorphism, and a morphism is an equivalence of a chain map, up to chain homotopy. This is a localization process [We94, §10.3] which has an $A_{\infty}$-version in [LO06].
@@
      The directed category $\mathbb{Z}$ indexes the diagram in $\mathcal{O}$ for each Lagrangian $L$, given by $\ldots\to\mathcal{Y}_{L^{k}}\xrightarrow[]{\mu^{2}(e_{L^{k}},-)}\mathcal{Y}_{L^{k+1}}\to\ldots$ for $k\in\mathbb{Z}$, with multiplication by quasi-units $e_{L^{k}}$. Define
      $\mathcal{Y}_{L^{\infty}}:=\underset{k\in\mathbb{Z}}{\text{hocolim}}(\mathcal{Y}_{L^{k}}).$
@@
      We will need the continuation maps of §3.2.4 to relate Floer complexes of different amounts of wrapping, allowing us to derive and localize the morphisms in §3.2.5.
@@
The homotopy direct limit over a collection of isotopies avoids the definition depending on the amount of wrapping. Note that in $\mathbb{C}^{*}$, using a quadratic Hamiltonian to wrap as in [fuk_intro, §4.2] is equivalent to a colimit by linear Hamiltonians, the original definition of the wrapped Fukaya category [AS10, §5.1], by [OT24, Proposition 6.1.9]. We will use this below in Subsection 3.3: Lagrangian gradings.
@@ 3.2.3. Quasi-isomorphisms: quasi-units
Following the convention of [AA24, Equation (3.42)], a quasi-unit roughly counts certain curves $\Sigma:=\mathbb{D}\backslash\{1\}\to Y^{0}$ with one moving Lagrangian boundary condition. The boundary condition is defined by $\rho^{-\tilde{\epsilon}g_{e}(s)}(L)$ for $g_{e}:\mathbb{R}\cong\partial\Sigma\to[k,k+1]$ monotonically increasing smooth and constant near the ends, meaning there is some $K\in\mathbb{R}$ so
@@
In particular, the Lagrangian isotopy is traversed in reverse along the boundary from $L^{k}$ to $L^{k+1}$.
The quasi-unit is the image of 1 under the PSS isomorphism $PSS:H^{*}(L)\to HF^{*}(L^{k},L^{k+1})$. We define it on the chain-level by a disc count,
(3.22) $e_{L^{k}}:=PSS(1)=\left[\sum_{p\in L^{k}\cap L^{k+1}}\sum_{\begin{subarray}{c}[u]:[u]=\beta\in\pi_{2}(Y^{0},\bigcup_{s\in\mathbb{R}}\rho^{-\tilde{\epsilon}g_{e}(s)}(L^{k}))\\ \text{ind}(u)=0\end{subarray}}\#\mathcal{M}(p;[u],J)e^{-2\pi\int_{\beta}\omega}\cdot p\in CF(L^{k},L^{k+1})\right],\\$